5. The pedals of a bicycle rotate in a circle with a diameter of 20 cm. What is the maximum torque a 58-kilogram rider can apply by putting all of her weight on one pedal?
Answer #1

The rider applies a force F = mg when she puts all the weight on the pedal.


Using the formula: torque = r*F, where r is the distance from the axis. (radius = d/2), we get:

torque = m*g*d/2 = 58*9.8*0.2/2 = 56.84 (N*m).